OpenStudy (anonymous):

find the exact solution of the following quadratic equation by using the quadratic formula A.{-7,1} D. {-1,7}

OpenStudy (lalaly):

whts the equation

OpenStudy (anonymous):

\[x^2-6x=7\]

OpenStudy (lalaly):

\[-b \pm \sqrt{b ^{2}-4ac } /2a\] \[6+\sqrt{36-4\times -7} /2\] = 7 and the same but with the minus so its D